OpenStudy (anonymous):

. The perimeter of a rectangle is 52. The length is 6 greater than the width. Find the length and width of this rectangle.

OpenStudy (anonymous):

2l + 2w = 52, so w + l = 26 so w + w + 6 = 26 so 2w = 20 so w = 10, l = 16